OpenStudy (anonymous):

2y + 4x = 18 2y - 3x = 4 A. (6, -3) B. (2, 5) C. (5, -1) D. (3, 3)

OpenStudy (dangerousjesse):

Solve the following system: {4 x+2 y = 18 | (equation 1) 2 y-3 x = 4 | (equation 2) Add 3/4 × (equation 1) to equation 2: {4 x+2 y = 18 | (equation 1) 0 x+(7 y)/2 = 35/2 | (equation 2) Divide equation 1 by 2: {2 x+y = 9 | (equation 1) 0 x+(7 y)/2 = 35/2 | (equation 2) Multiply equation 2 by 2/7: {2 x+y = 9 | (equation 1) 0 x+y = 5 | (equation 2) Subtract equation 2 from equation 1: {2 x+0 y = 4 | (equation 1) 0 x+y = 5 | (equation 2) Divide equation 1 by 2: {x+0 y = 2 | (equation 1) 0 x+y = 5 | (equation 2) Collect results: Answer: | | {x = 2 y = 5

OpenStudy (anonymous):

B?

OpenStudy (dangerousjesse):

Mhm ^-^

OpenStudy (anonymous):

thanks